# Sample size and power calculation for propensity score analysis of observational studies

### v2

@@ 3.1 Normal approximation of covariates
Assume $X_{1},X_{2},\dots$ is a sequence of independent random variables with mean 0 and variance 1. Let $\beta_{1},\beta_{2},\dots$ be a sequence of real numbers. If the following two conditions hold: (i) there exists a positive number $B$ such that $\mathbb{E}[X_{j}^{4}]<B$, and (ii) $\max_{1\leq j\leq p}\frac{\beta_{j}^{2}}{\sum_{j=1}^{p}\beta_{j}^{2}}=o\left(1\right)$, then
@@
The proof of Theorem 1 is relegated to Appendix A.2. Condition (i) ensures that the distribution of each covariate does not have heavy tails, and condition (ii) ensures that no single covariate dominates the linear combination. Though these two conditions may not always hold, empirical evidence from simulated and three real applications (see Appendix B.1) shows the normality generally holds in a wide range of scenarios.
@@ 3.2 Distribution of propensity scores
To determine the two parameters $\mu_{e}$ and $\sigma_{e}$ of the logit-normal distribution, we would need to impose constraints in the form of summary information of the treatment assignment. The first constraint, which is nearly always available, is the proportion of the treatment group in the intended study: $r=\Pr(Z_{i}=1)$. The second constraint is a measure, denoted as $\phi$, of the similarity between the covariate distributions in the treatment and control groups, $\Pr(X\mid Z=1)$ versus $\Pr(X\mid Z=0)$. There are many choices of the specification of $\phi$; important selection criteria include: (i) there should be 1-1 map between $(r,\phi)$ and $(\mu_{e},\sigma_{e})$ so that the data generating process of the propensity score is unique, and (ii) easy to compute the propensity score distribution (equivalently $\mu_{e},\sigma_{e}$) given $(r,\phi)$.
@@
The logit-normal distribution does not have a closed-form mean and thus it is generally difficult to solve for $(\mu_{e},\sigma_{e})$ given $(r,\phi)$. Instead, we propose to approximate the logit-normal distribution by a Beta distribution, which has a range on $[0,1]$ and is easy to compute. Applying the method in Aitchison and Shen, (1980), we can derive the “optimal” Beta function approximation of a logit-normal function by minimizing the Kullback-Leibler divergence between them. Specifically, for a logit-normal distribution $\mathcal{P}(\mu_{e},\sigma_{e}^{2})$ and its optimal approximated Beta distribution $\mathsf{Beta}(a,b)$, their parameters must satisfy the following relation: $\mu_{e}=\psi(a)-\psi(b)$ and $\sigma_{e}^{2}=\psi^{\prime}(a)+\psi^{\prime}(b)$, where $\psi$ and $\psi^{\prime}$ are the digamma and trigamma functions, respectively. Numerical examples in Appendix B.2 illustrate that this approximation is accurate for a wide range of logit-normal distributions.
If $e(X)\sim\mathsf{Beta}(a,b)$, then there is a closed-form relation between $a,b,r$:
$r=\frac{a}{a+b}.$ (9)
@@
Therefore, it suffices to identify the propensity score distribution if we find another easily computable relation between $a,b$ and the overlap measure $\phi$. The question now becomes how to specify such a $\phi$. Directly operating on the multivariate covariate $X$ is challenging; instead we operate on the propensity score $e(X)$, which is the coarsest balancing score (Rosenbaum and Rubin,, 1983). Let $f_{z}(u)$ be the density of the propensity score in arm $z$: $f_{z}(u)=\Pr(e(X)=u\mid Z=z)$. We propose to use the Bhattacharyya coefficient (Bhattacharyya,, 1943) to measure the covariate overlap between two groups:
$\phi\equiv\int_{0}^{1}\sqrt{f_{1}(u)f_{0}(u)}\,\mathrm{d}u.$ (10)
@@
The Bhattacharyya coefficient $\phi$ satisfies $0\leq\phi\leq 1$, with $\phi=0$ if and only if the support of $f_{0}$ and $f_{1}$ does not overlap, whereas $\phi=1$ if and only if $f_{0}=f_{1}$ almost everywhere. Therefore, $\phi$ is an appropriate measure of overlap; we term it as the overlap coefficient hereafter. We choose the Bhattacharyya coefficient because it has a major computable advantage in our context, established in the following proposition (the proof is relegated to Appendix A.3).
@@
$\phi=\frac{\Gamma(a+0.5)}{\sqrt{a}\Gamma(a)}\frac{\Gamma(b+0.5)}{\sqrt{b}\Gamma(b)}.$ (11)
Therefore, given $(r,\phi)$, we can obtain the Beta parameters $(a,b)$ by solving for Equation (9) and (11). Specifically, based on Equation (9), it is natural to reparameterize $(a,b)$ as $a=kr$ and $b=k(1-r)$, under which Equation (11) becomes
$\phi(k)=\frac{\Gamma(kr+0.5)}{\sqrt{kr}\Gamma(kr)}\frac{\Gamma(k(1-r)+0.5)}{\sqrt{k(1-r)}\Gamma(k(1-r))}.$ (12)
We write $\phi(k)$ to highlight that, given $r$, the overlap coefficient $\phi$ is a function of the unknown $k$. The following proposition establishes that, given $r$, there is a 1-1 map between $\phi$ and $k$, and equivalently, a 1-1 map between $(r,k)$ and $(a,b)$. The proof is relegated to Appendix A.4.
@@
The function, $\phi(k)$, defined in (12), is monotonically increasing in $k$ if $k>\max\left\{\frac{1}{2r},\frac{1}{2(1-r)}\right\}$.
Proposition 3 allows us to use the bisection method to solve for $k$ efficiently. Interestingly, our experience suggests that many other specification of the overlap coefficient do not lead to a 1-1 map to the propensity score distribution. In summary, as long as users specify $(r,\phi)$, we can solve for the Beta parameters $(a,b)$ in a computationally efficient fashion, based on which we further reverse the Beta approximation of $e(X)$ to obtain the logit-normal distribution parameters $(\mu_{e},\sigma_{e})$.
@@ 3.3 Distribution of potential outcomes
For simplicity, we first focus on continuous outcomes. The potential outcomes are usually specified as a linear model $Y(z)=\gamma_{z0}+X^{\prime}\gamma_{z}+\epsilon_{z}$, with $\mathbb{E}[\epsilon_{z}]=0$. Similarly as before, Theorem 1 motivates us to approximate the distribution of $X^{\prime}\gamma_{z}$ by a normally distributed random variable $W_{z}$.
@@
Recall that we approximated the distribution of the propensity score by $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, where $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$ is a normal random variable. Therefore, the correlation between the potential outcomes and the propensity score is captured by the correlation between $W_{e}$ and $W_{z}$. This motivates us to decompose $W_{z}$ into two parts, one linear in $W_{e}$ and the other orthogonal to $W_{e}$. Specifically, we can decompose $W_{z}=a_{z}W_{e}+W_{e}^{\perp}$ such that $\mathrm{cov}(W_{e},W_{e}^{\perp})=0$. Then, the potential outcomes can be expressed as $Y(z)=a_{z}W_{e}+W_{e}^{\perp}+\epsilon_{z}$, where both $W_{e}^{\perp}$ and $\epsilon_{z}$ are unknown normal random variables. Because both variables are independent of $W_{e}$, it is impossible to further identify $W_{e}^{\perp}$ from $\epsilon_{z}$; thus, with slight abuse of notation, we combine these two terms and denote the sum as $\epsilon_{z}$, under which the potential outcomes are modeled as
$Y(z)=a_{z}W_{e}+\epsilon_{z},\quad\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma^{2}_{z}).$ (13)
@@
Here we have six unknown parameters in total to characterize the potential outcome distributions: $a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2}$. In sample size calculation settings, users usually have prior knowledge of the mean and variance of the observed outcome in each treatment group (e.g. available in Table 1 of a previous study in medical papers), denoted as $E_{z}\equiv\mathbb{E}[Y\mid Z=z]$ and $S_{z}\equiv\operatorname{\mathbb{V}}[Y\mid Z=z]$, for $z=0,1$. These are different from their counterparts to the potential outcomes $\operatorname{\mathbb{E}}[Y(z)]$ and $\operatorname{\mathbb{V}}[Y(z)]$. In addition to these four quantities, we would need two more constraints to identify the six parameters; it is natural to use the correlation between the covariates and the observed outcomes, i.e., $R_{z}\equiv\mathrm{cor}[Y,W_{e}\mid Z=z]$. We postpone the detailed discussion of $R_{z}$ to Section 3.4.
@@
With the six specified values $(E_{1},E_{0},S_{1},S_{0},R_{1},R_{0})$, one can solve for the unknown parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ by the following equations (the proof is relegated to Appendix A.5)
$a_{z}$ $=R_{z}\sqrt{\frac{S_{z}}{\operatorname{\mathbb{V}}[W_{e}\mid Z=z]}},\quad\mu_{z}=E_{z}-a_{z}\mathbb{E}[W_{e}\mid Z=z],\quad\sigma_{z}^{2}=(1-R_{z}^{2}){S_{z}}.$ (14)
@@
Note that the conditional mean $\mathbb{E}[W_{e}\mid Z=z]$ and variance $\operatorname{\mathbb{V}}[W_{e}\mid Z=z]$ do not have closed forms. We propose to calculate them numerically by the following formula
@@
$\mathbb{E}[W_{e}\mid Z=z]$ $=\frac{\int_{\mathbb{R}}x\exp\left\{-\frac{1}{2\sigma_{e}^{2}}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}{\int_{\mathbb{R}}\exp\left\{-\frac{1}{2\sigma_{e}^{2}}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x},$ (15)
@@
$\operatorname{\mathbb{V}}[W_{e}\mid Z=z]$ $=\frac{\int_{\mathbb{R}}x^{2}\exp\left\{-\frac{1}{2\sigma_{e}^{2}}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}{\int_{\mathbb{R}}\exp\left\{-\frac{1}{2\sigma_{e}^{2}}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}-\left(\mathbb{E}[W_{e}\mid Z=z]\right)^{2}.$ (16)
@@ 3.4 Correlation between outcomes and covariates
The correlation parameter $R_{z}=\mathrm{cor}[Y,W_{e}\mid Z=z]$ may be challenging to interpret and specify. We first give an intuitive explanation. In observational studies, only the observed mean outcomes $\mathbb{E}[Y\mid Z=z]$, instead of the target potential outcomes $\mathbb{E}[Y(z)]$, are observable. If $R_{z}$ deviates far from zero, meaning that the covariates that are correlated with the propensity score are also correlated with the potential outcome, then $\mathbb{E}[Y(z)]$ would be more likely different from $\mathbb{E}[Y\mid Z=z]$ because the units in the opposite treatment group would likely have different covariates that lead to different potential outcomes. Therefore, as $R_{z}$ increases, larger sample sizes are needed to recover the counterfactual $\mathbb{E}[Y(z)]$ from the observed $\mathbb{E}[Y\mid Z=z]$.
To gain more technical insight into $R_{z}$, consider the simple setting in which all covariates are uncorrelated and standardized with mean 00 and variance $1$ within each treatment arm, i.e., $\operatorname{\mathbb{V}}[X\mid Z=z]=I_{p}$, where $I_{p}$ is an identity matrix of dimension $p$. Recall that the propensity score and the outcome models are $e(X)=\operatorname{\mathrm{e}xpit}(\beta_{0}+X^{\prime}\beta)$ and $Y(z)=\gamma_{z0}+X^{\prime}\gamma_{z}+\epsilon_{z}$, respectively. In this outcome model, the variance of $Y(z)$ can be decomposed into two components, $\mathbb{V}[X^{\prime}\gamma_{z}\mid Z=z]$ and $\mathbb{V}[\epsilon_{z}\mid Z=z]$, respectively. The first component originates from the variance of $X$ and can be explained by the individual covariates. The second component is the unexplainable variance introduced by $\epsilon_{z}$. Thus, these two components are often termed signal and noise, respectively. We define $\mathrm{SNR}_{z}=\mathbb{V}[X^{\prime}\gamma_{z}\mid Z=z]/\mathbb{V}[\epsilon_{z}\mid Z=z]$ as the signal-to-noise ratio of the outcome model. Then the correlation parameter
$R_{z}=\frac{\beta^{\prime}\gamma_{z}}{\sqrt{\beta^{\prime}\beta\cdot(\gamma_{z}^{\prime}\gamma_{z}+\mathbb{V}[\epsilon_{z}\mid Z=z])}}=\frac{\cos\langle\beta,\gamma_{z}\rangle}{\sqrt{1+\mathrm{SNR}_{z}^{-1}}},$ (17)
where the numerator $\cos\langle\beta,\gamma_{z}\rangle$ is the cosine of the angle between $\beta$ and $\gamma_{z}$, which measures the closeness in direction of the coefficients $\beta$ and $\gamma_{z}$. When the sets of covariates associated with the propensity scores and the potential outcomes, respectively, are distinct, $\beta$ and $\gamma_{z}$ are close to orthogonal, and thus $\cos\langle\beta,\gamma_{z}\rangle$ will be close to 1. This includes randomized experiments as a special case. In general cases, the angle can be obscure for researchers; nevertheless, the absolute value of $R_{z}$ is bounded by $1/\sqrt{1+\mathrm{SNR}_{z}^{-1}}$, or $\sqrt{\mathrm{SNR}_{z}}$ when $\mathrm{SNR}_{z}$ is small. The signal-to-noise ratio, $\mathrm{SNR}_{z}$, does not involve propensity scores; instead, it only measures how much variation in the outcomes can be explained by the covariates and is arguably easier to specify than $R_{z}$.
The expression of $R_{z}$ does not depend on the intercepts $\beta_{0}$ and $\gamma_{z0}$. Therefore, $R_{z}$ is invariant under transformation of the potential outcome models. Specifically, under homogeneous effects $Y(1)=Y(0)+\tau$, we always have $R_{1}=R_{0}$; under heterogeneous effects, simulations and real applications in Section 5 suggest that $R_{1}$ and $R_{0}$ are often similar. Thus, to simplify the implementation, we recommend to impose $R_{1}=R_{0}=R$ and treat $R$ as a sensitivity parameter. As shown in (19) and (24), the sample size formulas for both ATE and WATE depend only on the absolute value of $R$. Therefore, the range of $R$ can be reduced to $[0,\sqrt{\mathrm{SNR}_{z}}]$ instead of $[-1,1]$.

### v3

@@ 3.1 Normal approximation of linear combinations of covariates
Assume $X_{1},X_{2},\dots$ is a sequence of independent random variables with mean 0 and variance 1. Let $\beta_{1},\beta_{2},\dots$ be a sequence of real numbers. If the following two conditions hold: (i) there exists a positive number $B$ such that $\mathbb{E}(X_{j}^{4})<B$, and (ii) $\max_{1\leq j\leq p}{\beta_{j}^{2}}(\sum_{j=1}^{p}\beta_{j}^{2})^{-1}=o(1)$, then
@@
The proof of Theorem 1 is relegated to Supplementary Material A.2. Condition (i) ensures that the distribution of each covariate does not have heavy tails, and condition (ii) ensures that no single covariate dominates the linear combination. Though these two conditions may not always hold, empirical evidence from simulated and three real applications (see Supplementary Material B.1) shows the normality generally holds in a wide range of scenarios.
@@ 3.2 Distribution of propensity scores
To determine the two parameters $\mu_{e}$ and $\sigma_{e}$ of the logit-normal distribution, we would need to impose constraints in the form of summary information of the treatment assignment. The first constraint, which is usually available, is the proportion of the treatment group in the intended study: $r=\Pr(Z_{i}=1)$. The second constraint is a measure, denoted as $\phi$, of the similarity between the covariate distributions in the treatment and control groups, $\Pr(X\mid Z=1)$ versus $\Pr(X\mid Z=0)$. There are many choices of the specification of $\phi$; important selection criteria include: (i) there should be 1-1 map between $(r,\phi)$ and $(\mu_{e},\sigma_{e})$ so that the data generating process of the propensity score is unique, and (ii) easy to compute the propensity score distribution given $(r,\phi)$.
@@
The logit-normal distribution does not have a closed-form mean and thus it is generally difficult to solve for $(\mu_{e},\sigma_{e})$ given $(r,\phi)$. Instead, we propose to approximate the logit-normal distribution by a Beta distribution, which has a range on $[0,1]$ and is easy to compute. Applying the method in Aitchison and Shen, (1980), we can derive the “optimal” Beta function approximation of a logit-normal function by minimizing the Kullback-Leibler divergence between them. Specifically, for a logit-normal distribution $\mathcal{P}(\mu_{e},\sigma_{e}^{2})$ and its optimal approximated Beta distribution $\mathsf{Beta}(a,b)$, their parameters must satisfy the following relation: $\mu_{e}=\psi(a)-\psi(b)$ and $\sigma_{e}^{2}=\psi^{\prime}(a)+\psi^{\prime}(b)$, where $\psi$ and $\psi^{\prime}$ are the digamma and trigamma functions, respectively. Numerical examples in Supplementary Material B.2 illustrate that this approximation is accurate for a wide range of logit-normal distributions.
@@
If $e(X)\sim\mathsf{Beta}(a,b)$, then there is a closed-form relation between $a,b,r$: $r={a}/(a+b)$. Therefore, it suffices to identify the propensity score distribution if we find another easily computable relation between $a,b$ and the overlap measure $\phi$. The question now becomes how to specify such a $\phi$. Directly operating on the multivariate covariate $X$ is challenging; instead we operate on the propensity score $e(X)$, which is the coarsest balancing score (Rosenbaum and Rubin,, 1983). Let $f_{z}(u)$ be the density of the propensity score in arm $z$: $f_{z}(u)=\Pr(e(X)=u\mid Z=z)$. We propose to use the Bhattacharyya coefficient (Bhattacharyya,, 1943) to measure the covariate overlap between two groups:
$\phi\equiv\int_{0}^{1}\{f_{1}(u)f_{0}(u)\}^{1/2}\,\mathrm{d}u.$ (9)
@@
The Bhattacharyya coefficient $\phi$ satisfies $0\leq\phi\leq 1$, with $\phi=0$ if and only if the support of $f_{0}$ and $f_{1}$ does not overlap, whereas $\phi=1$ if and only if $f_{0}=f_{1}$ almost everywhere. Therefore, $\phi$ is an appropriate measure of overlap; we term it as the overlap coefficient hereafter. We choose the Bhattacharyya coefficient because it has a major computable advantage in our context, established in the following proposition (the proof is relegated to Supplementary Material A.3).
@@
$\phi=\frac{\Gamma(a+0.5)}{{a}^{1/2}\Gamma(a)}\frac{\Gamma(b+0.5)}{{b}^{1/2}\Gamma(b)}.$ (10)
Therefore, given $(r,\phi)$, we can obtain the Beta parameters $(a,b)$ by solving for Equations $r={a}/(a+b)$ and (10). Specifically, based on the former, we can reparameterize $(a,b)$ as $a=kr$ and $b=k(1-r)$, under which Equation (10) becomes
$\phi(k)=\frac{\Gamma(kr+0.5)}{(kr)^{1/2}\Gamma(kr)}\frac{\Gamma(k(1-r)+0.5)}{\{k(1-r)\}^{1/2}\Gamma(k(1-r))}.$ (11)
We write $\phi(k)$ to highlight that, given $r$, the overlap coefficient $\phi$ is a function of the unknown $k$. The following proposition establishes that, given $r$, there is a 1-1 map between $\phi$ and $k$, and equivalently, a 1-1 map between $(r,k)$ and $(a,b)$. The proof is relegated to Supplementary Material A.4.
@@
The function, $\phi(k)$, defined in (11), is monotonically increasing in $k$ if $k>\max\left\{(2r)^{-1},\{2(1-r)\}^{-1}\right\}$.
Proposition 2 allows us to use the bisection method to solve for $k$ efficiently. Interestingly, we found that many other specification of the overlap coefficient do not lead to a 1-1 map to the propensity score distribution. In summary, as long as users specify $(r,\phi)$, we can solve for the Beta parameters $(a,b)$ in a computationally efficient fashion, based on which we further reverse the Beta approximation of $e(X)$ to obtain the logit-normal distribution parameters $(\mu_{e},\sigma_{e})$.
@@ 3.3 Distribution of potential outcomes
For simplicity, we focus on continuous outcomes. The potential outcomes are usually specified as a linear model $Y(z)=\gamma_{z0}+X^{\prime}\gamma_{z}+\epsilon_{z}$, with $\mathbb{E}[\epsilon_{z}]=0$. Similarly as before, Theorem 1 motivates us to approximate the distribution of $X^{\prime}\gamma_{z}$ by a Gaussian random variable $W_{z}$.
@@
Recall that we approximated the distribution of the propensity score by $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, where $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$ is a normal random variable. Therefore, the correlation between the potential outcomes and the propensity score is captured by the correlation between $W_{e}$ and $W_{z}$. This motivates us to decompose $W_{z}$ into two parts, one linear in $W_{e}$ and the other orthogonal to $W_{e}$. Specifically, we can decompose $W_{z}=a_{z}W_{e}+W_{e}^{\perp}$ such that $\mathrm{cov}(W_{e},W_{e}^{\perp})=0$. Then, the potential outcomes can be expressed as $Y(z)=a_{z}W_{e}+W_{e}^{\perp}+\epsilon_{z}$, where both $W_{e}^{\perp}$ and $\epsilon_{z}$ are unknown normal random variables. Because both variables are independent of $W_{e}$, it is impossible to further identify $W_{e}^{\perp}$ from $\epsilon_{z}$; thus, with slight abuse of notation, we combine these two terms and denote the sum as $\epsilon_{z}$, under which the potential outcomes are modeled as $Y(z)=a_{z}W_{e}+\epsilon_{z},\quad\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma^{2}_{z}).$
@@
Here we have six unknown parameters in total to characterize the potential outcome distributions: $a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2}$. In sample size calculation settings, users usually have prior knowledge of the mean and variance of the observed outcome in each treatment group (e.g. available in Table 1 of a previous study in medical papers), denoted as $E_{z}\equiv\mathbb{E}(Y\mid Z=z)$ and $S_{z}^{2}\equiv\operatorname{\mathbb{V}}(Y\mid Z=z)$, for $z=0,1$. These are different from their counterparts to the potential outcomes $\operatorname{\mathbb{E}}\{Y(z)\}$ and $\operatorname{\mathbb{V}}\{Y(z)\}$. In addition to these four quantities, we would need two more constraints to identify the six parameters; it is natural to use the correlation between the covariates and the observed outcomes, i.e., $\rho_{z}\equiv\mathrm{cor}(Y,W_{e}\mid Z=z)$. We postpone the detailed discussion of $\rho_{z}$ to Section 3.4.
@@
With the six specified values $(E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1},\rho_{0})$, one can solve for the unknown parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ by the following equations (the proof is relegated to Supplementary Material A.5)
$a_{z}$ $=\rho_{z}|S_{z}|\{\operatorname{\mathbb{V}}(W_{e}\mid Z=z)\}^{-1},\quad\mu_{z}=E_{z}-a_{z}\mathbb{E}(W_{e}\mid Z=z),\quad\sigma_{z}^{2}=(1-\rho_{z}^{2}){S_{z}^{2}}.$ (12)
@@
The conditional mean $\mathbb{E}(W_{e}\mid Z=z)$ and variance $\operatorname{\mathbb{V}}(W_{e}\mid Z=z)$ do not have closed forms. We propose to calculate them numerically by the following formula
@@
$\mathbb{E}(W_{e}\mid Z=z)$ $=\frac{\int_{\mathbb{R}}x\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}{\int_{\mathbb{R}}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x},$ (13)
@@
$\operatorname{\mathbb{V}}(W_{e}\mid Z=z)$ $=\frac{\int_{\mathbb{R}}x^{2}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}{\int_{\mathbb{R}}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}-\left\{\mathbb{E}(W_{e}\mid Z=z)\right\}^{2}.$ (14)
@@ 3.4 Correlation between outcomes and covariates
We now interpret the seemingly opaque correlation parameter $\rho_{z}=\mathrm{cor}(Y,W_{e}\mid Z=z)$. Recall that the propensity score $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, thus $W_{e}$ summarizes the covariates that are predictive of the treatment assignment. Therefore, $\rho_{z}$ measures the confounder-outcome association: as $\rho_{z}$ increases, confounding covariates exert a greater influence on the outcome between treatments. The parameter $\rho_{z}$ is intrinsically related to a standard and interpretable quantity: the covariate-outcome association, $R_{z}=\mathrm{cor}(Y,\widetilde{Y}\mid Z=z)$, where $\widetilde{Y}$ is the component of $Y$ that can be explained by the covariates. Under the linear outcome model $Y(z)=X^{\prime}\gamma_{z}+\epsilon_{z}$, we have $\widetilde{Y}=X^{\prime}\gamma_{z}$ and the square of the covariate-outcome association, $R_{z}^{2}$, is equivalent to the R-squared statistic $R_{z,Y\mid X}=\operatorname{\mathbb{V}}(X^{\prime}\gamma_{z}\mid Z=z)/\operatorname{\mathbb{V}}(Y\mid Z=z)$. Then, we can derive the following expression to show that $\rho_{z}$ is bounded by $R_{z}$:
$\rho_{z}^{2}$ $=\left[\frac{\mathrm{cov}(Y,W_{e}\mid Z=z)}{\{\mathbb{V}(W_{e}\mid Z=z)\}^{1/2}\{\mathbb{V}(\widetilde{Y}\mid Z=z)\}^{1/2}}\right]^{2}{\frac{\mathbb{V}(\widetilde{Y}\mid Z=z)}{\mathbb{V}(Y\mid Z=z)}}=\mathrm{cor}^{2}(\widetilde{Y},W_{e}\mid Z=z)R_{z}^{2}\leq R^{2}_{z}.$ (15)
In randomized trials, covariates are independent of the treatment assignment by design and thus $\mathrm{cor}(\widetilde{Y},W_{e}\mid Z=z)=0$ and $\rho_{z}=0$. In observational studies, as the covariates associated with propensity scores differ more from the covariates associated with potential outcomes, $\mathrm{cor}(\widetilde{Y},W_{e}\mid Z=z)$ will get closer to 0. Equation (15) implies that users can bound $\rho_{z}^{2}$ by the $R_{z}^{2}$ by the R-squared value of a hypothetical regression of the outcome on the covariates, of which investigators often have an educated guess based on prior domain knowledge. Moreover, $\rho_{z}$ is invariant under linear transformation of the potential outcome models. Specifically, under homogeneous effects $Y(1)=Y(0)+\tau$, we always have $\rho_{1}=\rho_{0}$; under heterogeneous effects, simulations and real examples in Section 4 suggest that $\rho_{1}$ and $\rho_{0}$ are usually similar. Also, as will be shown in formula (17), the minimal sample size for $\tau^{\mbox{\tiny{ATE}}}$ depends on $\rho$ only through $\rho^{2}$. Therefore, in practice, we recommend to impose $\rho_{1}=\rho_{0}=\rho$ for simplicity, and conduct a sensitivity analysis of $|\rho|$ in the range $(0,|R_{Y|X}|)$ with a domain knowledge informed specification of $R_{Y|X}$, or set $\rho=0$ when such information is unavailable, which could lead to a mild underestimate of the sample size.